OpenStudy (anonymous):

Igor who is 4ft 8inches tall wishes to find the height of an oak tree in front of his castle. He walks along the shadow of the tree until the end of his shadow excatly overlaps the end of the tree's shadow. At that point he is 22ft from the foot of the tree and 11ft from the end of the shadows. How tall is the oak tree? help would be amazing!!<33

OpenStudy (anonymous):

What do you not understand?

OpenStudy (anonymous):

I could give you the answer but I'd like to help you learn how to solve these questions so you won't need help again.

OpenStudy (anonymous):

Well I don't understand how to draw the picture, the word probblems are weird to me

OpenStudy (anonymous):

Ok I will draw and attach sketches step by step.

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

No problem, so first it says Igor is 4 feet and 8 inches tall. So draw a line from the "ground" and label its measure as 4 feet and 8 inches.

OpenStudy (anonymous):

He wants to find the height of the oak tree, we'll make another line to the ground and label it X because we don't know what it is yet.

OpenStudy (anonymous):

mkay yeah

OpenStudy (anonymous):

gotcha what next?

OpenStudy (anonymous):

So it would be 4.8/11=x/11?

OpenStudy (anonymous):

Whoops I misread something! I apologize! And no it's not 4.8 because it's 8 out of 12 inches not 8 out of 10.

OpenStudy (anonymous):

So what would the equation be?

OpenStudy (anonymous):

Ok first you understand the the height of the tree will be similar to the height of the man?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok the sketch will actually look like the following and I will explain:

OpenStudy (anonymous):

He follows the shadow until the end of his shadow is at the same point as the end of the tree's shadow. In this process he has moved 22 feet in front of the tree.

OpenStudy (anonymous):

From where he stands, the shadows end 11 feet away from him.

OpenStudy (anonymous):

Now, see the 2 similar triangles?

OpenStudy (anonymous):

Tell me what you think the ratio of similarity should be.

OpenStudy (anonymous):

maybe 56/132=x/132?

OpenStudy (anonymous):

Good guess but no. If you look closely, you can see that to find the bottom side measurement of the larger triangle you must add 11 feet + 22 feet.

OpenStudy (anonymous):

So the tree's shadow is actually 33 feet long.

OpenStudy (anonymous):

and the man's is 11 feet long.

OpenStudy (anonymous):

So, in inches the similarity should be: 56/132 = x/396 Make sense?

OpenStudy (anonymous):

oh okay I see!:) thanks so much!!!

OpenStudy (anonymous):

I'm glad I could help :)

OpenStudy (anonymous):

Don't forget to convert the answer back to feet after you've finished!

OpenStudy (anonymous):

oh i totally forgot thanks!